OpenStudy (anonymous):

what property would 2*5*y=y*5*z be? 1. symmetric 2. transitive 3. associative property of mult. 4. communitive property of mult.

OpenStudy (anonymous):

none

OpenStudy (anonymous):

2*5*y=y*5*2 is the number 4th

OpenStudy (anonymous):

so communitive property of multipication?

OpenStudy (anonymous):

i assumed you put a 'z' instead of 2

Directrix (directrix):

Should 2*5*y=y*5*z be 2*5*y=y*5*2 ? Typo?

OpenStudy (anonymous):

Oh, true :b

OpenStudy (anonymous):

umm let me check

OpenStudy (anonymous):

yes sorry ;P

OpenStudy (anonymous):

what about 2(x+6) - (x+1) = 16?

OpenStudy (anonymous):

^x=5/2

OpenStudy (anonymous):

a. 3 b. 5 c. 2 d.1

OpenStudy (anonymous):

you shouldn't be asking question within a question..post it

OpenStudy (anonymous):

ok? haha sorry